# Adventures in Demand Analysis Using AI

### v1

@@ Abstract
Keywords: Causal Inference; Deep Embeddings; Causal Fine-Tuning; Text; Images; Difference-in-Difference; Panel Data; Debiased Machine Learning; Demand Analysis; Foundation Models.
@@ 1 Introduction
Almost a century ago, The Journal of the American Statistical Association published early empirical studies of demand that applied statistical methods to measure how consumers respond to price changes: Work by scholars such as Wright (1929), Working (1943), Schultz (1933), Mills (1931, 1937a, 1937b), and Stigler (1939) moved economics from theory towards quantitative measurement. By doing so, this work provided a foundation for econometrics, a field of statistical analysis focusing on economic problems. Their research established a tradition of using data to understand market behavior and inform economic models.
@@
Our approach contributes to multiple strands of the literature. It extends empirical demand analysis by employing AI-generated, multimodal representations of products. Our work also builds on the emerging intersection of econometrics and machine learning (Athey and Imbens, 2019; Mullainathan and Spiess, 2017; Varian, 2014; Chernozhukov et al., 2018a) and complements recent studies that apply AI-based text analysis and other modern methods to economic questions (Belloni et al., 2014; Bajari et al., 2023; Compiani et al., 2023). In doing so, it provides a framework for combining flexible product representations with established econometric tools for identification and inference. It also introduces the idea that embeddings can and should be fine-tuned with causal inference in mind—which is critical in our context and potentially useful in other applications.
Our key empirical result is that AI-based embeddings are strong determinants/modifiers of the price elasticity function. Interestingly, we also find that these embeddings are not major confounders of the relationship between price and quantity. Thus, while standard homogeneous price-effect models can yield ballpark estimates of average elasticity, they can severely understate or overstate the price elasticity for certain sets of products. These findings constitute important new empirical insights.
@@ 2.1 The Data and Measurement of Prices and Quantities
Our analysis uses a data set of toy cars from Amazon.com, compiled and provided by the data aggregator Keepa.com. For each item $i$, we collected its sales rank and price at time points spanning from April to December 2023. We also gathered each product’s description, image, and additional tabular features (e.g., its subcategory on Amazon.com), as summarized in Table 1. Figure 1 illustrates a typical product page containing the product image and description. Overall, our data set comprises $N=9{,}613$ unique products.
@@
Each period, indexed by $t=1,\ldots,T$, spans 4 weeks. We have $T=9$ periods in total, each separated by 1 week. We structure our data set this way to limit inter-temporal feedback in our price elasticity analysis in Section 3. We also examine temporal changes in these signals, $\Delta Q_{it}:=Q_{it}-Q_{i(t-1)}$ and $\Delta P_{it}:=P_{it}-P_{i(t-1)}.$
@@
Price Lowest offered price for a new product, excluding shipping and handling fees
@@
Buy box Binary; 1 if the product delivery is fulfilled by Amazon
@@
Our use of the time-averaged inverse sales rank is an approximation motivated by modeling the latent, true quantities $Q^{*}_{it}$ as independent draws from an underlying distribution at each time $t$. If an item has sales rank $k$, the quantity sold should be distributed as the $k^{\text{th}}$ order statistic from a sample of size $N$. In the case of Pareto distributions with shape $\vartheta$, a unit change in $\log Q^{*}_{it}$, $\Delta\log Q^{*}_{it}$, corresponds on average to a change of $-\vartheta\cdot\Delta\log Q^{*}_{it}$ in the log sales rank, $-Q_{it}$. This motivates our use of the sales rank. Using this approximation, He and Hollenbeck (2020) estimate $\hat{\vartheta}\approx 0.5$ for toys on Amazon.com; consequently, we can multiply our estimates by $1/\hat{\vartheta}\approx 2$ to obtain rough estimates of the price elasticity of demand. More generally, we can quantify the connection between our estimates and price elasticities under various assumptions about the distribution of sales (Office of National Statistics, 2020).
@@ 2.2 Using AI to Represent Products
To convert product data into useful numerical features, we employ various encoding models based upon the transformer neural network architecture proposed by Vaswani et al. (2017). We convert text descriptions into embeddings $T_{i}$ using language models such as RoBERTa (Liu et al., 2019) or LLaMA 3 (Touvron et al., 2023), convert images into dense embeddings $I_{i}$ using the BEiT model (Bao et al., 2022), and transform tabular data into embeddings using the SAINT model (Somepalli et al., 2022). The Appendix provides additional implementation details.
@@
[…]
@@
[…]
@@ 2.3 Evaluating the Embeddings
After obtaining the embeddings, we must assess whether they effectively represent the products and “understand” their characteristics. We first take the concatenated embeddings $E_{i}=\bigl{(}T_{i},I_{i}\bigr{)}$, where $T_{i}$ also includes tabular embeddings, and then apply a Johnson–Lindenstrauss projection of these embeddings onto a 256-dimensional vector $\,\bar{E}_{i}$; Johnson (1984). This projection approximately preserves distances and is therefore considered (at least approximately) information-lossless. We then center and normalize the embeddings so they lie on a hypersphere:
@@ 2.3.1 Qualitative Assessment
Figure 3: 3d representation of product embeddings (with image) and five clusters
Figure 4: 3d representation of product embeddings (no image) and five clusters
While mainly illustrative, these visuals suggest that text-only embeddings lie in a lower-dimensional space, missing valuable image-based information absent from text descriptions. This observation is supported by Tables 2 and 4, which show that clusters formed from text+image embeddings have more visually coherent centroids and exhibit greater internal homogeneity—confirming the importance of multimodal data.
@@
Cluster 0 Premium-Detail Replicas of Emergency, Service, and Racing Vehicles for Realistic Play [Uncaptioned image]
Cluster 1 Oversized and Interactive Truck Playsets Offering Transforming Functions and Action-Packed Fun [Uncaptioned image]
Cluster 2 Pull-Back, Friction-Powered, and Easily Transforming Vehicles for Fast, Engaging Adventures [Uncaptioned image]
Cluster 3 A Variety of Diecast Classics and Specialty Vehicles for Collectors and Everyday Play [Uncaptioned image]
Cluster 4 Iconic Movie-Inspired 1:55 Scale Diecast Cars Perfect for Storytelling and Roleplay [Uncaptioned image]
@@ 2.3.2 Quantitative Assessment
As shown in Table 5, simple linear regressions using only tabular data perform poorly. Boosted trees yield substantial gains in predictive accuracy, and neural networks with text embeddings perform even better. Including image embeddings offers further improvements, though the additional gains are modest (e.g., a 1.5 percentage-point increase in $R^{2}$ for $Q_{it}$). Nonetheless, these gains are meaningful.
We also assess each model’s ability to predict changes in quantities and prices, rather than their levels. As expected, predicting changes is notably more difficult, leading to a sharp decline in performance. Even the best models achieve only about 18% $R^{2}$ for changes in quantity and 1% $R^{2}$ for changes in price. Despite this drop, the results underscore that AI-generated features provide strong predictive power for quantity levels and some improvement—albeit smaller—for predicting changes.
@@
  • Principal Components (PCA):
    $X^{pc}_{i,k}:=\gamma_{k}^{T}X^{e}_{i},\quad X^{pc}_{i}:=(X^{pc}_{i,k})_{k=1}^{K},$
    where $\gamma_{k}$ is the $k$-th eigenvector of the covariance matrix of $X^{e}_{i}$, corresponding to the $k$-th largest eigenvalue.
  • Centroid Similarities (CS):
    $X^{sim}_{i,k}:=c_{k}^{T}X^{e}_{i},\quad X^{sim}_{i}:=(X^{sim}_{i,k})_{k=1}^{K},$
    where $c_{k}$ is the centroid (mean) of the $k$-th cluster identified by $k$-means.
These vectors capture how similar the embedding vectors are to principal axes of variation—either principal components or $k$-means cluster centroids—using cosine similarity. (E.g., since ${c}_{k}$ and ${X}^{e}_{i}$ lie on the unit hypersphere, their inner product ${c}_{k}^{\top}{X}^{e}_{i}$ directly represents the cosine similarity). As shown in Table 6, using only five principal components or five centroid similarities can retain nearly all the prediction-relevant information contained in the original embeddings. In particular, when using a boosted tree, these compressed embeddings nearly match the performance of a deep neural network that uses the entire text and image inputs. This finding can simplify downstream tasks. In this paper, we rely primarily on centroid similarities, as they appear more interpretable than principal components for our application.
@@
Linear Reg [all tabular] 22.08% 17.10% 6.11% 0.36%
Boosted Trees Reg [all tabular] 45.84% 20.37% 13.21% -2.60%
Deep Learning Reg [text only; invariant tabular] 50.99% 63.12% 0.01% -0.00%
Deep Learning Reg [image and text; invariant tabular] 51.24% 64.67% 0.05% -0.07%
Deep Learning Reg [text only; all tabular] 60.04% 65.15% 18.41% 0.76%
Deep Learning Reg [image and text; all tabular] 61.77% 65.43% 11.35% 1.03%
Note: All models are trained on a training set and scores are evaluated on a test set. Predictions use lagged values of time-varying controls.
@@
Linear Reg [+5 PCAs] 48.69% 59.07% 6.15% 0.40%
Linear Reg [+ 5 Similarities] 48.60% 57.70% 6.16% 0.43%
Linear Reg [+256 Embeddings] 51.08% 59.66% 5.81% -0.04%
Boosted Trees Reg [+5 PCAs] 56.02% 63.17% 14.99% -1.28%
Boosted Trees Reg [+5 Similarities] 56.12% 62.28% 15.37% -0.30%
Boosted Trees Reg [+256 Embeddings] 55.73% 64.05% 14.03% -0.72%
Note: All models are trained on a training set and scores are evaluated on a test set. Predictions use lagged values of time-varying controls.
@@ 3.1 Initial Approach and Challenges
where $g_{t}(\cdot)$ is a function describing how the control variables influence the outcome over time. We allow $g_{t}$ to vary with $t$. The notation $\mathrm{L}[Y\mid P,X]$ denotes the projection of the random variable $Y$ onto the space of partially linear prediction rules of the form $a\,P+g(X)$.
In practice, directly estimating this model often suggests a very small price sensitivity (or “elasticity”), captured by the coefficient $\delta$: $\delta\approx[0,-0.2]$. From a causal perspective, this result is implausible: the notion that a price change exerts virtually no effect on ranking or sales is counterintuitive, as it implies that raising prices would only marginally reduce the quantity sold. In other words, this finding suggests that certain key confounders—such as latent quality or visibility—are not adequately controlled for, resulting in a strongly biased estimate.
@@ 3.2 The Causal Dynamic Model and Regressions
To address the issue above, we introduce a simple dynamic panel data model to guide our statistical analysis. We can view the outcomes and key variables as arising from the following structural equation model (SEM):
@@
This specification defines an autoregressive model in which the quantity signal $Q_{it}$ depends on the price signal $P_{it}$ and other state variables $S_{it}$. The state variables include lagged quantity and price, $Q_{i,t-1}$ and $P_{i,t-1}$, time-invariant product characteristics $X_{i}$ (captured through embeddings), and time-varying characteristics $X_{it}^{o}$ (such as ratings and the number of reviews). Among these variables, the lagged quantity $Q_{i,t-1}$ is arguably a key confounder, reflecting both product visibility and quality—a conclusion reinforced by our empirical findings below. In particular, including the lagged quantity in the model substantially shifts the estimated price elasticity into a more plausible range.
@@
Identification of both the CACE and the ACE in this setting follows from computing the conditional expectation of $Q_{it}$ while conditioning on $P_{it}$ (the treatment) and $S_{it}$ (the observed confounders). Indeed, conditioning on $S_{it}$ blocks non-causal sources of association between the outcome and the treatment (Pearl, 1995). Including further lags of state variables $S_{it}$ is not strictly necessary under the Markovian structure, but it can serve as a useful specification check. Below, we also discuss potential threats from unobserved confounders, such as time-varying demand shocks that may affect both the outcome and the treatment.
@@
where $\gamma_{t}(S_{it})\;:=\;\mathrm{E}\bigl{[}q_{t}\bigl{(}S_{it},\varepsilon_{it}\bigr{)}\bigr{]},$ so the CACE function $\alpha_{t}(S_{it})$ appears as the heterogeneous slope in (4). The ACE parameter $\alpha_{t}$ then follows by averaging the CACE function over $S_{it}$.

### v2

@@ Abstract
Keywords: Causal Inference; Deep Embeddings; Causal Fine-Tuning; Text; Images; Panel Data; Debiased Machine Learning; Demand Analysis; Foundation Models.
@@ 1 Introduction
Almost a century ago, The Journal of the American Statistical Association published early empirical studies of demand that applied statistical methods to measure how consumers respond to price changes. Work by scholars such as Wright (1929), Working (1943), Schultz (1933), Mills (1931, 1937a, 1937b), and Stigler (1939) moved economics from theory towards quantitative measurement. By doing so, this work provided a foundation for econometrics, a field of statistical analysis focusing on economic problems. Their research established a tradition of using data to understand market behavior and inform economic models.
@@
Our approach contributes to multiple strands of the literature. It extends empirical demand analysis by employing AI-generated, multimodal representations of products. Our work also builds on the emerging intersection of econometrics and machine learning (Belloni et al., 2014; Varian, 2014; Athey and Imbens, 2019; Mullainathan and Spiess, 2017; Chernozhukov et al., 2018a) and complements recent studies that apply AI-based text analysis and other modern methods to economic questions (Bajari et al., 2023; Compiani et al., 2023). In doing so, it provides a framework for combining flexible product representations with established econometric tools for identification and inference. It also introduces the idea that embeddings can and should be fine-tuned with causal inference in mind—which is critical in our context and potentially useful in other applications.
Our main empirical result is that AI-based embeddings are important effect modifiers: they capture substantial heterogeneity in price elasticity across products. Interestingly, our findings also suggest that these embeddings are not major confounders of the price-quantity relationship. Although highly predictive of price and quantity levels, they are weak predictors of the temporal changes we use to identify elasticities. Thus, while standard homogeneous price-effect models can yield reasonable estimates of the average elasticity, they can severely understate or overstate elasticity for certain sets of products. These findings constitute important new empirical insights.
@@ 2.1 The Data and Measurement of Prices and Quantities
Our analysis uses a data set of toy cars from Amazon.com, compiled and provided by the data aggregator Keepa.com. For each item $i$, we collected its sales rank and price at time points spanning from March 2023 to January 2024. We also gathered each product’s description, image, and additional tabular features (e.g., its subcategory on Amazon.com), as summarized in Table 1. Figure 1 illustrates a typical product page containing the product image and description. Overall, our data set comprises N=7,226N=7{,}226 unique products.
@@
Each period, indexed by $t=1,\ldots,T$, spans 4 weeks. We have $T=12$ periods in total, which are directly adjacent to each other. We structure our data set this way to limit inter-temporal feedback in our price elasticity analysis in Section 3. We also examine temporal changes in these signals, $\Delta Q_{it}:=Q_{it}-Q_{i(t-1)}$ and $\Delta P_{it}:=P_{it}-P_{i(t-1)}.$
@@
Price Buy Box price for a new product, excluding shipping and handling fees
@@
Buy box is FBA Binary; 1 if the product delivery is fulfilled by Amazon
@@
Our use of the time-averaged inverse sales rank is an approximation motivated by modeling the latent, true quantities $Q^{*}_{it}$ as independent draws from an underlying distribution at each time $t$. If an item has sales rank $k$ at time $t$, the quantity sold should be distributed as the $(n-k+1)^{\text{th}}$ order statistic from a sample of size $n$, denoted $Q_{(n-k+1)t}^{*}$. In the case of Pareto distributions with shape $\vartheta$, i.e. $\Pr(Q^{*}_{it}>t)=C_{\vartheta}\cdot t^{-\vartheta}$, we have the asymptotic approximation $\log\mathrm{E}[Q_{(n-k+1)t}]\sim C_{\vartheta,n}-\vartheta^{-1}\log(k).$ [^1] Moments of order statistics for the Pareto distribution are given explicitly by Malik (1966) (see also Vännman, 1976); combining the result with Stirling’s approximation gives $\mathrm{E}[Q_{(n-k+1)t}^{*}]\sim C_{\vartheta}(k/n)^{-1/\vartheta}$ as $n\to\infty$. This motivates our use of the negative log sales rank, $Q_{it}$. Using this approximation, He and Hollenbeck (2020) estimate $\hat{\vartheta}\approx 0.5$ for toys on Amazon.com; consequently, we can multiply our estimates by $1/\hat{\vartheta}\approx 2$ to obtain rough estimates of the price elasticity of demand. More generally, one can quantify the connection between our estimates and price elasticities under various assumptions about the distribution of sales (Office for National Statistics, 2020).
@@ 2.2 Using AI to Represent Products
To convert product data into useful numerical features, we employ various encoding models based upon the transformer neural network architecture proposed by Vaswani et al. (2017). We convert text descriptions into embeddings $T_{i}$ using language models such as RoBERTa (Liu et al., 2019) or LLaMA (Touvron et al., 2023), convert images into dense embeddings $I_{i}$ using the BEiT model (Bao et al., 2022), and transform tabular data into embeddings using the SAINT model (Somepalli et al., 2022). The Online Appendix provides additional implementation details.
@@
Consider the example sentence: $S=\text{``Well made diecast model truck with metal body.''}$ We create a masked version: $W=\text{``Well made [m] model truck with [m] body''}.$ The original sequence $S$ serves as an auxiliary label that the model attempts to reconstruct from the corrupted input $W$. By applying this approach to billions of sentences, models learn to capture syntactic and semantic relationships without explicitly annotated labels.
@@
$\quad A^{tx}_{i}$$X^{in}_{i}=\begin{bmatrix}\mathrm{Text}_{i}\\ \mathrm{Image}_{i}\end{bmatrix}$$E_{i}\coloneqq\begin{bmatrix}T_{i}\\ I_{i}\end{bmatrix}$$\{\hat{Q}_{it},\hat{P}_{it}\}_{t=1}^{T}.$$\quad A^{im}_{i}$$\scriptstyle{e}$$\scriptstyle{m}$
@@ 2.3 Evaluating the Embeddings
After obtaining the embeddings, we must assess whether they effectively represent the products and “understand” their characteristics. We first take the concatenated embeddings $E_{i}=\bigl(T_{i},I_{i}\bigr)$, where $T_{i}$ also includes tabular embeddings, and then apply a Johnson–Lindenstrauss projection of these embeddings onto a 256-dimensional vector $\bar{E}_{i}$; Johnson and Lindenstrauss (1984). This projection approximately preserves distances and is therefore considered (at least approximately) information-preserving. We then center and normalize the embeddings so they lie on a hypersphere:
@@ 2.3.1 Qualitative Assessment
Figure 3: 3D-representation of product embeddings (with image) and five clusters
Figure 4: 3D-representation of product embeddings (no image) and five clusters
While mainly illustrative, these visuals suggest that text-only embeddings lie in a lower-dimensional space, missing valuable image-based information absent from text descriptions. This observation is supported by Tables 2 and 4, which show that clusters formed from text+image embeddings have more visually coherent centroids and greater internal homogeneity—confirming the importance of multimodal data.
@@
Cluster 0 Movie-Themed Character Collectibles and Customization Kits [Uncaptioned image]
Cluster 1 Realistic Large-Scale Utility Trucks with Functional Components [Uncaptioned image]
Cluster 2 1:64 Scale Die-Cast Real-World Automotive Replicas [Uncaptioned image]
Cluster 3 Medium-Scale Pull-Back Models and Wooden Train Sets [Uncaptioned image]
Cluster 4 Interactive Stunt Vehicles, Launchers, and Adventure Playsets [Uncaptioned image]
@@ 2.3.2 Quantitative Assessment
As shown in Table 5, simple linear regressions using only tabular data perform poorly. Boosted trees yield substantial gains in predictive accuracy, and neural networks with text embeddings perform even better. Including image embeddings offers further improvements, though the additional gains are modest. Nonetheless, these gains are meaningful.
We also assess each model’s ability to predict changes in quantities and prices, rather than their levels. As expected, predicting changes is notably more difficult, leading to a sharp decline in performance. Even the best models achieve only about 15% $R^{2}$ for changes in quantity and 1.5% $R^{2}$ for changes in price. Despite this drop, the results underscore that AI-generated features provide strong predictive power for quantity levels and some improvement—albeit smaller—for predicting changes.
@@
  • Principal Components (PCA): $X^{pc}_{i,k}:=\gamma_{k}^{T}X^{e}_{i},\ X^{pc}_{i}:=(X^{pc}_{i,k})_{k=1}^{K},$ where $\gamma_{k}$ is the $k$-th eigenvector of the covariance matrix of $X^{e}_{i}$, corresponding to the $k$-th largest eigenvalue.
  • Centroid Similarities (CS): $X^{sim}_{i,k}:=c_{k}^{T}X^{e}_{i},\ X^{sim}_{i}:=(X^{sim}_{i,k})_{k=1}^{K},$ where $c_{k}$ is the centroid (mean) of the $k$-th cluster identified by $k$-means.
These vectors capture how similar the embedding vectors are to principal axes of variation—either principal components or $k$-means cluster centroids—using cosine similarity. As shown in Table 6, using only five principal components or five centroid similarities can retain nearly all the prediction-relevant information contained in the original embeddings. In particular, when using a boosted tree, these compressed embeddings nearly match the performance of a deep neural network that uses the entire text and image inputs. This finding can simplify downstream tasks. In this paper, we rely primarily on centroid similarities, as they appear more interpretable than principal components for our application.
@@
Linear Reg [all tabular] 20.84% 15.82% 7.08% 0.64%
Boosted Trees Reg [all tabular] 47.51% 18.54% 15.13% 0.00%
Deep Learning Reg [text only; invariant tabular] 50.31% 65.74% 9.48% 1.24%
Deep Learning Reg [image and text; invariant tabular] 50.44% 67.26% 10.89% 1.15%
Deep Learning Reg [text only; all tabular] 60.65% 64.89% 14.48% 1.01%
Deep Learning Reg [image and text; all tabular] 59.53% 66.69% 12.17% 1.65%
Note: All models are trained on a training set and scores are evaluated on a test set. Predictions use lagged values of time-varying controls. Negative $R^{2}$ values are set to 0.
@@
Linear Reg [+5 PCAs] 46.47% 60.63% 7.06% 0.64%
Linear Reg [+ 5 Similarities] 47.27% 57.02% 7.07% 0.64%
Linear Reg [+256 Embeddings] 51.24% 62.14% 5.47% 0.00%
Boosted Trees Reg [+5 PCAs] 53.22% 64.36% 13.79% 0.00%
Boosted Trees Reg [+5 Similarities] 52.58% 62.53% 14.63% 0.00%
Boosted Trees Reg [+256 Embeddings] 55.28% 65.64% 10.11% 0.00%
Note: All models are trained on a training set and scores are evaluated on a test set. Predictions use lagged values of time-varying controls. Negative $R^{2}$ values are set to 0.
@@ 3.1 Initial Approach and Challenges
where $g_{t}(\cdot)$ is a function describing how the control variables influence the outcome over time. We allow $g_{t}$ to vary with $t$. The notation $\mathrm{L}[Y\mid P,X]$ denotes the projection of the random variable $Y$ onto the space of partially linear prediction rules of the form aP+g(X)a\,P+g(X).
In practice, directly estimating this model often suggests a very small price sensitivity (or “elasticity”), captured by the coefficient $\delta\in[-0.3,\,-0.1]$. From a causal perspective, this result is implausible: the notion that a price change exerts virtually no effect on ranking or sales is counterintuitive, as it implies that raising prices would only marginally reduce the quantity sold. In other words, this finding suggests that certain key confounders—such as latent quality or visibility—are not adequately controlled for, resulting in a strongly biased estimate.
@@ 3.2 The Causal Dynamic Model and Regressions
To address the issue above, we introduce a simple dynamic panel data model to guide our statistical analysis. We can view the outcomes and key variables as arising from the following structural equation model (SEM).
Assumption 1.
(Structural Equation Model)
  1. a)
    It holds
@@
  2. b)
    The variables $Q_{it}$, $P_{it}$, and $S_{it}$ are square-integrable.
This specification defines an autoregressive model in which the quantity signal $Q_{it}$ depends on the price signal $P_{it}$ and other state variables $S_{it}$. The state variables include lagged quantity and price, $Q_{i,t-1}$ and $P_{i,t-1}$, time-invariant product characteristics $X_{i}$ (captured through embeddings), and time-varying characteristics $X_{it}^{o}$ (such as ratings and the number of reviews). Among these variables, the lagged quantity $Q_{i,t-1}$ is arguably a key confounder, reflecting both product visibility and quality—a conclusion reinforced by our empirical findings below. In particular, including the lagged quantity substantially shifts the estimated price elasticity into a more plausible range.
@@
Identification of both the CACE and the ACE in this setting follows from computing the conditional expectation of $Q_{it}$ while conditioning on $P_{it}$ (the treatment) and $S_{it}$ (the observed confounders). Indeed, conditioning on $S_{it}$ blocks non-causal sources of association between the outcome and the treatment (Pearl, 1995). Including further lags of state variables $S_{it}$ is not strictly necessary under the Markovian structure, but it can serve as a useful specification check. Such identification assumes the absence of potential unobserved confounders, such as time-varying demand shocks that may affect both the outcome and the treatment. We further discuss limitations of our identification strategy and possible alternative strategies in Section 3.6, below.
@@
where $\gamma_{t}(S_{it})\;:=\;\mathrm{E}\bigl[q_{t}\bigl(S_{it},\varepsilon_{it}\bigr){\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}{|\,S_{it}}}\bigr],$ so the CACE function $\alpha_{t}(S_{it})$ appears as the heterogeneous slope in (4). The ACE parameter $\alpha_{t}$ then follows by averaging the CACE function over $S_{it}$.